Concordance Invariance of

Levine-Tristram Signatures of Links

Matthias Nagel, Mark Powell

Received: September 7, 2016 Revised: December 14, 2016

Communicated by Ursula Hamenstädt

Abstract. We determine for which complex numbers on the unit circle the Levine-Tristram signature and the nullity give rise to link concordance invariants.

2010 Mathematics Subject Classification: primary: 57M25, secondary: 57M27, 57N70 Keywords and Phrases: link concordance, Levine-Tristram signatures, nullity

1. Introduction

Let L ⊂ S^{3} be an m-component oriented link in the 3-sphere. Each connected, oriented Seifert surface F for L has a bilinear Seifert form defined by

V : H_{1}(F; ℤ) × H_{1}(F; ℤ) → ℤ
(p[x], q[y]) ↦ pq lk(x^{−}, y),

where p, q ∈ ℤ, x, y are simple closed curves on F with associated homology classes [x], [y], and x^{−} is a push-off of x in the negative normal direction of F. Given a unit modulus complex number z ∈ S^{1}∖{1}, choose a basis for H_{1}(F; ℤ) and define the hermitian matrix

B(z) := (1 − z)V + (1 − \overline{z})V^{T} .

The Levine-Tristram signature σ_{L}(z) of L at z is defined to be the signature of B(z), namely the number of positive eigenvalues minus the number of negative eigenvalues. The nullity η_{L}(z) of L at z is the dimension of the null space of B(z). Both quantities can be shown to be invariants of the S-equivalence class of the Seifert matrix, and are therefore link invariants [Lev69, Tri69].

We say that two oriented m-component links L and J are concordant if there is a flat embedding into S^{3} × I of a disjoint union of m annuli A ⊂ S^{3} × I, such that the oriented boundary of A satisfies

∂A = −L ⊔ J ⊂ −S^{3} ⊔ S^{3} = ∂(S^{3} × I).

An m-component link L is slice if it is concordant to the m-component unlink. The purpose of this paper is to answer the following question: for which values of z are σ_{L}(z) and η_{L}(z) link concordance invariants? We work in the topological category, in order to obtain the strongest possible results. In order to state our main theorem, we need one more definition.

Definition 1.1. A complex number z ∈ S^{1}∖{1} is a Knotennullstelle if there exists a Laurent polynomial p(t) ∈ ℤ[t, t^{−1}] with p(1) = ±1 and p(z) = 0.

Note that a complex number z ∈ S^{1}∖{1} is a Knotennullstelle if and only if there exists a knot K whose Alexander polynomial Δ_{K} has the property that Δ_{K}(z) = 0. This follows from the fact that all Laurent polynomials q ∈ ℤ[t, t^{−1}] with q(1) = ±1 and q(t) = q(t^{−1}) can be realised as Alexander polynomials of knots [BZ03, Theorem 8.13]. Here is our main theorem.

Theorem 1.2. The link invariants σ_{L}(z) and η_{L}(z) are concordance invariants if and only if z ∈ S^{1}∖{1} does not arise as a Knotennullstelle.

Discussion of previously known results. The first point to note is that, due to J. C. Cha and C. Livingston [CL04], when z is a Knotennullstelle neither σ_{L}(z) nor η_{L}(z) are link concordance invariants.

Theorem 1.3 (Cha, Livingston). For any Knotennullstelle z ∈ S^{1}∖{1}, there exists a slice knot K with σ_{K}(z) ≠ 0 and η_{K}(z) ≠ 0.

Given a polynomial p(t) with p(1) = ±1 and p(z) = 0, Cha and Livingston construct a matrix V with V − V^{T} nonsingular, with det(tV − V^{T}) equal to p(t)p(t^{−1}), such that the upper left half-size block contains only zeroes, and such that σ(B(z)) ≠ 0. Such a matrix can easily be realised as the Seifert matrix of a slice knot. Some positive results on concordance invariance are also known. For z a prime power root of unity, σ_{L}(z) and η_{L}(z) are concordance invariants; see [Mur65], [Tri69] and [Kau78]. D. Cimasoni and V. Florens [CF08] dealt with multivariable signature and nullity concordance invariants, but again only at prime power roots of unity. For the signature and nullity at algebraic numbers away from prime power roots of unity, we could not find any statements or results in the literature pertaining to our question. Levine [Lev07] studied the question in terms of ρ-invariants, but only discussed concordance invariance away from the roots of the Alexander polynomial. By changing the rules slightly, one can obtain a concordance invariant for all z. The usual method is to define a function that is the average of the two one-sided limits of the Levine-Tristram signature function. Let z = e^{iθ} ∈ S^{1}, and consider:

\overline{σ}_{L}(z) := \frac{1}{2}( lim_{ω→θ_{+}} σ(B(e^{iω})) + lim_{ω→θ_{−}} σ(B(e^{iω}))).

Since prime power roots of unity are dense in S^{1}, this averaged signature function yields a concordance invariant at every z ∈ S^{1}. The earliest explicit observation of this that we could find was by Gordon in the survey article [Gor78]. One can also consider the averaged nullity function, to which similar remarks apply:

\overline{η}_{L}(z) := \frac{1}{2}( lim_{ω→θ_{+}} η(B(e^{iω})) + lim_{ω→θ_{−}} η(B(e^{iω}))).

In particular this is also a link concordance invariant. Note that the function σ_{L}: S^{1}∖{1} → ℤ is continuous away from roots of the Alexander polynomial det(tV − V^{T}) of L. More generally one can consider the torsion Alexander polynomial Δ_{L}^{Tor} of L, which by definition is the greatest common divisor of the (n − r) × (n − r) minors of tV − V^{T}, where n is the size of V and r is the minimal nonnegative integer for which the set of minors contains a nonzero polynomial. The function σ_{L} is continuous away from the roots of the torsion Alexander polynomial Δ_{L}^{Tor}, by [GL15, Theorem 2.1] (their A_{L} is our Δ_{L}^{Tor}). Thus if z is not a root of the torsion Alexander polynomial of any link, the signature cannot jump at that value, and the signature function σ_{L}(z) equals the averaged signature function \overline{σ}_{L}(z) there. Since the averaged function is known to be a concordance invariant, the non-averaged function is also an invariant when z is not the root of any link’s Alexander polynomial. The excitement happens when z is the root of the Alexander polynomial of some link, but is not the root of an Alexander polynomial of any knot. The averaged and non-averaged signature functions can differ at such z, but nevertheless both are concordance invariants. In Section 2 we will give an example which illustrates this difference, and gives an instance where the non-averaged function is more powerful. Similar examples were given in [GL15], but only with jumps occurring at prime power roots of unity. Finally we remark that our proof of Theorem 1.2 covers the previously known cases of prime power roots of unity and transcendental numbers, as well as the new cases.

Organisation of the paper. The rest of the paper is organised as follows. In Section 2, we give an example of two links that are not concordant, where we use the signature and nullity functions at a root of their Alexander polynomials, which is not a prime power root of unity, to detect this fact. Section 4 proves that the nullity is a concordance invariant, and the corresponding fact for signatures is proven in Sections 5 and 6.

Acknowledgements. We thank Stefan Friedl, Pat Gilmer, Chuck Livingston and Andrew Ranicki for helpful discussions. In particular [GL15] inspired the question that led to this paper. We also thank the referee for helpful feedback. M.

Powell is supported by an NSERC Discovery grant. M. Nagel is supported by a CIRGET Postdoctoral Fellowship.

2. An application

In the introduction, for a link L we defined the signature function σ_{L}(z) and the nullity function η_{L}(z), for each z ∈ S^{1}∖{1}. From the characterisation in Theorem 1.2, one easily finds new values z for which it was not previously known that σ(z) and η(z) are concordance invariants. In Proposition 2.3, by exhibiting the obligatory explicit example, we show that these values give obstructions to concordance that are independent from previously known obstructions coming from the signature and nullity functions. We finish the section by constructing, in Proposition 2.5, a family of such examples for any algebraic number on S^{1}. Before the construction, we collect some facts on the set of roots of Alexander polynomials of links. We say that a complex number z ∈ S^{1}∖{1} is a Linknullstelle if z is a root of a non-vanishing single variable Alexander polynomial of some link. We have the following inclusions:

{Knotennullstellen} ⊂ {Linknullstellen} ⊂ S^{1}∖{1}
∪
\{ \begin{matrix}prime power \\ roots of 1 \end{matrix} \}

We will see that these inclusions are strict. The two subsets of the set of Linknullstellen are disjoint, since no prime power root of unity can be a root of a polynomial that augments to ±1, because the corresponding cyclotomic polynomial augments to the prime. Moreover, the union of the Knotennullstellen and the prime power roots of unity is not exhaustive.

Lemma 2.1.

(1) The set of Linknullstellen coincides with the set of algebraic numbers in S^{1}∖{1}.

(2) The number z_{0} = \frac{3+4i}{5} ∈ S^{1} is an algebraic number, which is neither a Knotennullstelle nor a root of unity.

Proof. Let z ∈ S^{1}∖{1} be an algebraic number, so that p(z) = 0 for some p ∈ ℤ[t]. Let

q(t) := (t − 1)^{3}p(t)p(t^{−1}) ∈ ℤ[t, t^{−1}].

We claim that there is a link L with single variable Alexander polynomial Δ_{L}(t) = q(t). Choose a 2-variable polynomial P (x, y) ∈ ℤ[x^{±1}, y^{±1}] with P (t, t) = p(t). Let

Q(x, y) := (x − 1)(y − 1)P (x, y)P (x^{−1}, y^{−1}).

A corollary [Hil12, Corollary 8.4.1] to Bailey’s theorem [Bai77] states that any polynomial Q(x, y) in ℤ[x^{±1}, y^{±1}], with Q = \overline{Q} up to multiplication by ±x^{k}y^{ℓ}, and such that (x − 1)(y − 1) divides Q, is the Alexander polynomial of some 2-component link of linking number zero. Thus there exists a 2-component link L with 2-variable Alexander polynomial Q(x, y).

The single variable Alexander polynomial Δ_{L}(t) is obtained from the 2-variable Alexander polynomial of a 2-component link Q(x, y) as (t − 1)Q(t, t) [BZ03, Remark 9.18]. But

(t − 1)Q(t, t) = (t − 1)^{3}P (t, t)P (t^{−1}, t^{−1}) = (t − 1)^{3}p(t)p(t^{−1}) = q(t).

This completes the proof of the claim and therefore of (1): the set of Linknullstellen is the set of algebraic numbers lying on S^{1}∖{1}. For (2), first observe that the complex number z_{0} := \frac{3+4i}{5} has unit modulus and that z_{0} is a zero of the polynomial

p(t) := 5t^{2} − 6t + 5,

and therefore is an algebraic number. Note that no cyclotomic polynomial divides the polynomial p(t). This can be checked for the first six by hand, and the rest have degree larger than 2. From Abel’s irreducibility theorem, we learn that z_{0} is not a zero of a cyclotomic polynomial and thus is not a root of unity. Since p(1) = 4 and p(t) is irreducible over ℤ[t], z_{0} is not the root of any polynomial that augments to ±1. As a result, z_{0} is not a Knotennullstelle. □

Next we describe links L and L′ whose signature and nullity functions are equal everywhere on S^{1}∖{1} apart from at z_{0}, which will be a root of the Alexander polynomials of L and L′. We find these links by realising suitable Seifert forms.

Example 2.2. Consider the following Seifert matrix:

V := (\begin{matrix}0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 \\ 0 & 0 & 0 & 0 & 0 & 5 & −4 & 4 \\ 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 \\ 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 \\ 1 & −1 & 0 & 0 & 0 & 0 & 0 & 0 \\ 0 & 5 & −1 & 0 & 0 & 0 & 0 & 0 \\ 0 & −4 & 1 & −1 & 0 & 0 & 0 & 0 \\ 0 & 4 & 0 & 1 & 0 & 0 & 0 & 1 \end{matrix}) .

This matrix represents the Seifert form of the 3-component link L given by the boundary of the Seifert surface shown in Figure 1. As usual, a box with n ∈ ℤ inside denotes n full right-handed twists between two bands, made without introducing any twists into the individual bands. To see what we mean, observe that there are three instances in the figure of one full left-handed twist, otherwise known as −1 full right-handed twists. The left-most twist is between the bands labelled e_{1} and e_{5}. To obtain the Seifert matrix, note that the beginning of each of the eight bands is labelled e_{i}, for i = 1, …, 8. Orient the bands clockwise and compute using V_{ij} = lk(e_{i}^{−}, e_{j}), where the picture is understood to show the positive side of the Seifert surface. Produce a link L′ from L by removing the single twist in the right-most band, labelled e_{8} in Figure 1. This gives rise to a Seifert matrix V′ for L′ which is the same as V, except that the bottom right entry is a 0 instead of a 1.

\begin{matrix}e_{1} \\ e_{2} \\ e_{5} \\ e_{3} \\ e_{6} \\ e_{4} \\ e_{7} \\ e_{8} \\ −5 \\ 4 \\ −4 \end{matrix}

Figure 1. Realisation of the Seifert form V.

Consider the sesquilinear form B over ℚ[t^{±1}] determined by the matrix

(1 − t)V + (1 − t^{−1})V^{T} .

The form B splits into a direct sum of sesquilinear forms. For a Laurent polynomial p(t) ∈ ℚ[t^{±1}], abbreviate the form given by the 2 × 2 matrix

(\begin{matrix}0 & p(t) \\ p(t^{−1}) & 0 \end{matrix}) .

by [p(t)]. A calculation shows that B is congruent to the form

[t − 1] ⊕ [t − 1] ⊕ [t − 1] ⊕ (\begin{matrix}0 & q(t) \\ q(t^{−1}) & −t^{−1} + 2 − t \end{matrix}) ,

where the polynomial q(t) is

q(t) = t^{−1} · (t − 1)^{3} · (5t^{2} − 6t + 5).

On the other hand the corresponding sesquilinear form B′ over ℚ[t^{±1}] for L′ is equivalent to

[t − 1] ⊕ [t − 1] ⊕ [t − 1] ⊕ [q(t)].

Proposition 2.3. Let z_{0} denote the algebraic number \frac{3+4i}{5}. The links L and L′ constructed in Example 2.2 have the following properties.

(1) If z is a root of unity, then σ_{L}(z) = σ_{L′}(z) and η_{L}(z) = η_{L′}(z).

(2) The averaged signature and nullity functions agree, i.e.

\overline{σ}_{L}(z) = \overline{σ}_{L′}(z) and \overline{η}_{L}(z) = \overline{η}_{L′}(z)

for all z ∈ S^{1}∖{1}.

(3) The signatures and nullities of L and L′ at z_{0} differ:

σ_{L}(z_{0}) ≠ σ_{L′}(z_{0}) and η_{L}(z_{0}) ≠ η_{L′}(z_{0}),

and so L is not concordant to L′.

Proof. Note that for any z ∈ ℂ∖{0, 1} with q(z) ≠ 0, the form B(z) over ℂ is nonsingular and metabolic. The same holds for B′(z). This implies that the signatures sign B(z) and sign B′(z) vanish. The nullities η_{L}(z), η_{L′}(z) are also both zero. Since the roots of q(z) are exactly z_{0} and \overline{z_{0}}, which are not roots of unity by Lemma 2.1, we obtain the first statement of the proposition. We also see that the averaged signature function on S^{1}∖{1} and the averaged nullity function are identically zero, so we obtain the second statement. From Lemma 2.1, we know that z_{0} := \frac{3+4i}{5} is not a Knotennullstelle, and

sign B(z_{0}) = sign (\begin{matrix}0 & 0 \\ 0 & \frac{4}{5} \end{matrix}) = 1.

Thus σ_{L}(z_{0}) = 1 = η_{L}(z_{0}). On the other hand, for L′ the matrix B′(z_{0}) is a 2 × 2 zero matrix, so we have that σ_{L′}(z_{0}) = 0 and η_{L′}(z_{0}) = 2. Both signatures and the nullities at z_{0} differ, so L and L′ are not concordant by Theorem 1.2. □

Remark 2.4. One can also see that L and L′ are not concordant using linking numbers.

A more systematic study of the construction of the example above leads to the following proposition.

Proposition 2.5. Let q(t) ∈ ℤ[t] be a polynomial. Then there exists a natural number k > 0 and a link L with Alexander polynomial Δ_{L}(t) =^{.} q(t^{−1})q(t)(t − 1)^{k} up to units in ℤ[t, t^{−1}] such that

(1) the form B(z) of L is metabolic and nonsingular for all z ∈ S^{1}∖{1} which are not roots of q(t), so σ_{L}(z) = 0.

(2) if z_{0} ≠ 1 is a root of q(t) of unit modulus, then σ_{L}(z_{0}) ≠ 0.

The proof of this proposition is based on ideas from [CL04].

Proof. Consider the size n + 1 square matrix P with entries in ℤ[y] given by

P (y) := (\begin{matrix}1 & y & 0 & & & ya_{1} \\ 0 & 1 & y & & & ya_{2} \\ ⋮ & & ... & ... & & ⋮ \\ & & & 1 & y & ya_{n−1} \\ 0 & & & 0 & 1 & ya_{n} \\ y & 0 & … & 0 & 0 & 0 \end{matrix}) ,

with a_{i} integers. Over ℤ[y^{±1}], the matrix P can be transformed via invertible row operations and column operations to the matrix

A(y) = (\begin{matrix}1 & 0 & 0 & & & p(y) \\ 0 & 1 & 0 & & & 0 \\ ⋮ & & ... & ... & & ⋮ \\ & & & 1 & 0 & 0 \\ 0 & & & 0 & 1 & 0 \\ y & 0 & … & 0 & 0 & 0 \end{matrix})

with p(y) = b_{1}(y) where b_{k}(y) ∈ ℤ[y] is defined by the recursion b_{k−1}(y) := y · (a_{k} − b_{k}(y)) and b_{n}(y) := y · a_{n}. Notice that, up to units, we can arrange p(y) to be any polynomial in ℤ[y^{±1}] by choosing n sufficiently large and then suitable entries a_{k} ∈ ℤ. That is, multiply by y^{ℓ} so that the lowest order term is the linear term, and take (−1)^{i}a_{i} to be the coefficient of y^{i−1} in p(y), for i = 2, …, n + 1. Pick the entries a_{k} so that if we evaluate p(y) at (t−1) we get the equality p(t− 1) = q(t)(t − 1)^{k} for a suitable integer k. Now consider the block matrix

V := (\begin{matrix}0 & V^{u} \\ V^{b} & Q(1) \end{matrix})

with

V^{u} = (\begin{matrix}0 & 1 & 0 & & & a_{1} \\ 0 & 0 & 1 & & & a_{2} \\ ⋮ & & ... & ... & & ⋮ \\ & & & 0 & 1 & a_{n−1} \\ 0 & & & 0 & 0 & a_{n} \\ 1 & 0 & … & 0 & 0 & 0 \end{matrix}) V^{b} = (\begin{matrix}−1 & 0 & 0 & & & 1 \\ 1 & −1 & 0 & & & 0 \\ 0 & 1 & ... & ... & & ⋮ \\ ⋮ & & ... & −1 & 0 & 0 \\ 0 & & & 1 & −1 & 0 \\ a_{1} & a_{2} & … & a_{n−1} & a_{n} & 0 \end{matrix})

and

Q(y) = (\begin{matrix}0 & … & 0 & 0 \\ ⋮ & ... & ⋮ & ⋮ \\ 0 & … & 0 & 0 \\ 0 & … & 0 & y \end{matrix}) .

The matrix V is the Seifert matrix of a link as V − V^{T} is the intersection form of a genus n oriented surface with three boundary components. Let L be such a link, necessarily a 3-component link. We remark in passing that the matrix V from Example 2.2 is not a special case of the matrix V defined in the current proof, although it is close to being so. Recall that B(z) = (1 − z)V + (1 − \overline{z})V^{T} = (\overline{z} − 1) · (zV − V^{T} ). The matrix V was constructed in such a way that

B(z) = (\begin{matrix}0 & (\overline{z} − 1) · P (z − 1) \\ (z − 1) · P^{T} (\overline{z} − 1) & Q(−\overline{z} − z + 2) \end{matrix}) .

Using the transformations associated to the above row and column operations, we see that B(z) is congruent to

B(z) ∼ (\begin{matrix}0 & (\overline{z} − 1) · A(z − 1) \\ (z − 1) · A^{T} (\overline{z} − 1) & Q(−\overline{z} − z + 2) \end{matrix}) .

Note that the matrix Q is unchanged by this congruency, because in the corresponding sequence of row and column operations, it never happens that the last row or column is added to another row or column. We complete the proof of the proposition by showing that indeed the link L has the required properties. If z ∈ S^{1}∖{1} is not a zero of q(t), then also p(z) ≠ 0. Consequently, the form B(z) is nonsingular and metabolic. On the other hand, if z ∈ S^{1}∖{1} is a root of q(t), then also p(z) = 0. In this case the Levine-Tristram form B(z) is a sum

B(z) = M ⊕ (\begin{matrix}0 & 0 \\ 0 & −\overline{z} − z + 2 \end{matrix})

with M nonsingular and metabolic. Thus σ_{L}(z) = 1. □

Remark 2.6. Replace Q(1) with Q(0) in the construction of the matrix V in the proof of Proposition 2.5, to obtain a matrix V′. Using the same construction as in Example 2.2, the matrices V and V′ give rise to links L and L′ respectively, such that

η_{L}(z) = η_{L′}(z) and σ_{L}(z) = σ_{L′}(z)

for every z ∈ S^{1} that is not a root of q(t). Analogously to Example 2.2, L and L′ are not concordant, but again this can also be seen using linking numbers. This leads to the following question. Does there exist a pair of links L and L′, with the same pairwise linking numbers, whose signature and nullity functions can only tell the concordance classes of the links apart at an isolated algebraic numbers z, z̄ ∈ S^{1} that are roots of the Alexander polynomial Δ_{L} = Δ_{L′}.

3. Twisted homology and integral homology isomorphisms

Now we begin working towards the proof of Theorem 1.2. Fix z ∈ S^{1}∖{1} to be a unit complex number that is not the root of any polynomial p(t) ∈ ℤ[t] with p(1) = ±1 i.e. z is not a Knotennullstelle. We denote the classifying space for the integers ℤ by Bℤ, which has the homotopy type of the circle S^{1}. Given a CW complex X, a map X → Bℤ induces a homomorphism π_{1}(X) → ℤ. This determines a representation

α: ℤ[π_{1}(X)] → ℤ[ℤ] ⟶^{ev_{z}} ℂ

of the group ring of the fundamental group of X, with respect to which we can consider the twisted homology

H_{i}(X; ℂ^{α}) := H_{i} (ℂ ⊗_{ℤ[π_{1}(X)]} C_{∗}(X̃)) .

Let Σ ⊂ ℤ[ℤ] be the multiplicative subset of polynomials that map to ±1 under the augmentation ε: ℤ[ℤ] → ℤ, that is Σ = {p(t) ∈ ℤ[ℤ] | |p(1)| = 1}.

By inverting this subset we obtain the localisation Σ^{−1}ℤ[ℤ] of the Laurent polynomial ring. This has the following properties.

(i) The canonical map ℤ[ℤ] → Σ^{−1}ℤ[ℤ] is an inclusion, since ℤ[ℤ] is an integral domain.

(ii) For any ℤ[ℤ]-module morphism f: M → N of finitely generated free ℤ[ℤ]-modules such that the augmentation

ε(f) = Id ⊗f: ℤ ⊗_{ℤ[ℤ]} M → ℤ ⊗_{ℤ[ℤ]} N

is an isomorphism, we have that

Id ⊗f: Σ^{−1}ℤ[ℤ] ⊗_{ℤ[ℤ]} M → Σ^{−1}ℤ[ℤ] ⊗_{ℤ[ℤ]} N

is also an isomorphism. The second property can be reduced to the following. Assume A is a matrix over ℤ[ℤ] such that ε(A) is invertible. Consequently, we have det(ε(A)) = ±1 and as ε(det(A)) = det(ε(A)), we deduce that det(A) ∈ Σ. Therefore, the determinant det(A) is invertible in the localisation Σ^{−1}ℤ[ℤ] and so is the matrix A over Σ^{−1}ℤ[ℤ]. As the unit modulus complex number z that we have fixed is not a Knotennullstelle, the representation α defined above factors through the localisation,

i.e. evaluation at z determines a ring homomorphism Σ^{−1}ℤ[ℤ] ⟶^{Σ^{−1} ev_{z}} ℂ such that the ring homomorphisms ℤ[ℤ] ⟶^{ev_{z}} ℂ and

ℤ[ℤ] → Σ^{−1}ℤ[ℤ] ⟶^{Σ^{−1} ev_{z}} ℂ

coincide.

Lemma 3.1. Let f: X → Y be a map of finite CW complexes over S^{1}, that is there are maps g: X → S^{1} and h: Y → S^{1} such that h ∘ f = g, and suppose that

f_{∗}: H_{i}(X; ℤ) ⟶^{≅} H_{i}(Y; ℤ)

is an isomorphism for all i. Then

f_{∗}: H_{i}(X; ℂ^{α}) ⟶^{≅} H_{i}(Y; ℂ^{α})

is also an isomorphism for all i.

The lemma follows [COT03, Proposition 2.10]. The difference is that we use the well-known refinement that one does not need to invert all nonzero elements. We give the proof for the convenience of the reader. This is adapted from the proof given in [FP12].

Proof. The algebraic mapping cone D_{∗} := 𝒞 (f_{∗}: C_{∗}(X; ℤ) → C_{∗}(Y; ℤ)) has vanishing homology, and comprises finitely generated free ℤ-modules. Therefore it is chain contractible. We claim that the chain contraction can be lifted to a chain contraction for 𝒞 (f_{∗}: C_{∗}(X; Σ^{−1}ℤ[ℤ]) → C_{∗}(Y; Σ^{−1}ℤ[ℤ])), the mapping cone over the localisation Σ^{−1}ℤ[ℤ]. To see this, let s: D_{∗} → D_{∗+1} be a chain contraction, that is we have that ∂s_{i} +s_{i−1}∂ = Id_{D_{i}} for each i. Define D̃_{∗} := 𝒞 (f_{∗}: C_{∗}(X; ℤ[ℤ]) → C_{∗}(Y; ℤ[ℤ])) and consider ε: D̃_{∗} → D_{∗} = ℤ ⊗_{ℤ[ℤ]} D̃_{∗}, induced by the augmentation map. Denote E_{∗} := 𝒞 (f_{∗}: C_{∗}(X; Σ^{−1}ℤ[ℤ]) → C_{∗}(Y; Σ^{−1}ℤ[ℤ])) and note that there is an inclusion D̃_{i} → E_{i} = Σ^{−1}ℤ[ℤ] ⊗_{ℤ[ℤ]} D̃_{i}, induced by the localisation. Lift s to a map s̃: D̃_{∗} → D̃_{∗+1}, as in the next diagram

\xymatrix{D̃_{∗} \ar@{-->}[r]^{s̃} \ar[d]^{ε} & D̃_{∗+1} \ar[d]^{ε} \\ D_{∗} \ar[r]^{s} & D_{∗+1}.}

The lifts exist since all modules are free and ε is surjective. But then we have that

f := ds̃+ s̃d: D̃_{∗} → D̃_{∗}

is a morphism of free ℤ[ℤ]-modules whose augmentation ε(f) is an isomorphism. Thus by property (ii) of Σ^{−1}ℤ[ℤ], f is also an isomorphism over Σ^{−1}ℤ[ℤ], and so s̃ determines a chain contraction for E_{∗}. We therefore have that E_{∗} = C_{∗}(Y, X; Σ^{−1}ℤ[ℤ]) ≃ 0 as claimed. Next, tensor E_{∗} with ℂ over the representation α, to get that

ℂ^{α} ⊗_{Σ^{−1}ℤ[ℤ]} C_{∗}(Y, X; Σ^{−1}ℤ[ℤ]) = C_{∗}(Y, X; ℂ^{α}) ≃ 0.

Thus H_{i}(Y, X; ℂ^{α}) = 0 for all i and so f_{∗}: H_{i}(X; ℂ^{α}) ⟶^{≅} H_{i}(Y; ℂ^{α}) is an isomorphism for all i as desired. □

4. Concordance invariance of the nullity

In this section we show concordance invariance of the nullity function away from the set of Knotennullstellen.

Definition 4.1 (Homology cobordism). A cobordism (W^{n+1}; M^{n}, N^{n}) between n-manifolds M and N is said to be a ℤ-homology cobordism if the inclusion induced maps H_{i}(M; ℤ) → H_{i}(W; ℤ) and H_{i}(N; ℤ) → H_{i}(W; ℤ) are isomorphisms for all i ∈ ℤ.

Theorem 4.2. Suppose that oriented m-component links L and J are concordant and that z ∈ S^{1}∖{1} is not a Knotennullstelle. Then η_{L}(z) = η_{J}(z).

Proof. As in the statement suppose that z ∈ S^{1}∖{1} is not a Knotennullstelle. Denote the exterior of the link L by X_{L} := S^{3}∖νL. As above, let V be a matrix representing the Seifert form of L with respect to a Seifert surface F and a basis for H_{1}(F; ℤ). We assert that the matrix zV − V^{T} presents the homology H_{1}(X_{L}; ℂ^{α}). This can be seen as follows. Consider the infinite cyclic cover \overline{X}_{L} corresponding to the kernel of the homomorphism π_{1}(X_{L}) → ℤ, defined as the composition of the abelianisation π_{1}(X_{L}) → H_{1}(X_{L}; ℤ) ≅ ℤ^{m}, followed by the map (x_{1}, …, x_{m}) ↦ ∑_{i=1}^{m} x_{i} i.e. each oriented meridian is sent to 1 ∈ ℤ. A decomposition of \overline{X}_{L} and the associated Mayer-Vietoris sequence [Lic97, Theorem 6.5] give rise the following presentation

ℂ[t^{±1}] ⊗_{ℂ} H_{1}(F; ℂ) ⟶^{tV −V^{T}} ℂ[t^{±1}] ⊗_{ℂ} H_{1}(F; ℂ)^{∨} → H_{1}(\overline{X}_{L}; ℂ) → 0,

where H_{1}(F; ℂ)^{∨} is the dual module Hom_{ℂ}(H_{1}(F; ℂ), ℂ). Apply the rightexact functor ℂ^{α}⊗_{ℂ[t^{±1}]} to this sequence, to obtain the sequence

ℂ^{α} ⊗_{ℂ} H_{1}(F; ℂ) ⟶^{zV −V^{T}} ℂ^{α} ⊗_{ℂ} H_{1}(F; ℂ)^{∨} → ℂ^{α} ⊗_{ℂ[t^{±1}]} H_{1}(\overline{X}_{L}; ℂ) → 0.

As H_{0}(\overline{X}_{L}; ℂ) ≅ ℂ, we have that Tor_{1}^{ℂ[t^{±1}]}(H_{0}(\overline{X}_{L}; ℂ), ℂ^{α}) = 0 by the projective resolution

0 → ℂ[t^{±1}] →^{·(1−t)} ℂ[t^{±1}] → ℂ → 0

and z ≠ 1. Since ℂ[t^{±1}] is a principal ideal domain, we can apply the universal coefficient theorem for homology to deduce that ℂ^{α} ⊗_{ℂ[t^{±1}]} H_{1}(\overline{X}_{L}; ℂ) = H_{1}(X_{L}; ℂ^{α}). This completes the proof of the assertion that zV − V^{T} presents the homology H_{1}(X_{L}; ℂ^{α}). Next observe that (\overline{z} − 1)(zV − V^{T}) = (1 − z)V + (1 − \overline{z})V^{T} presents the same module as zV − V^{T}, since \overline{z} − 1 is nonzero. The dimension of H_{1}(X_{L}; ℂ^{α}) therefore coincides with the nullity η_{L}(z), which is by definition the nullity of the matrix (1 − z)V + (1 − \overline{z})V^{T}. Now, let A ⊂ S^{3} × I be a union of annuli giving a concordance between L and J, and let W := S^{3} × I∖νA. Then W is a ℤ-homology bordism between X_{L} and X_{J}; this is a straightforward computation with Mayer-Vietoris sequences or with Alexander duality; see for example [FP14, Lemma 2.4]. Thus by two applications of Lemma 3.1, with Y = W and X = X_{L} and X = X_{J} respectively, we see that H_{1}(X_{L}; ℂ^{α}) ≅ H_{1}(W; ℂ^{α}) ≅ H_{1}(X_{J}; ℂ^{α}), and so the nullities of L and J agree. We need that z is not a Knotennullstelle in order to apply Lemma 3.1. □

5. Identification of the signature with the signature of a 4-manifold

In the proof of Theorem 4.2, a key step was to reexpress the nullity η(z) of the form B(z) as a topological invariant of a 3-manifold, and then to use the bordism constructed from a concordance to relate the invariants. An analogous approach is used here to obtain the corresponding statement for the signature. Everything in this section is independent of whether z is a Knotennullstelle. Recall that we fixed an oriented m-component link L ⊂ S^{3}, and that we picked a connected Seifert surface F for L. Denote the link complement by X_{L} := S^{3}∖νL. First note that the fundamental class [F] ∈ H_{2}(F, ∂F; ℤ) of the Seifert surface F is independent of the choice of F. This follows from the fact that its Poincaré dual is characterised as the unique cohomology class ξ ∈ H^{1}(X_{L}; ℤ) mapping each meridian μ to ξ(μ) = 1. The boundary of F ⊂ S^{3}∖νL is a collection of embedded curves in the boundary tori that we refer to as the attaching curves. The attaching curves together with the meridians determine a framing of each boundary torus of X_{L}. Also, this framing depends solely on [F], since the connecting homomorphism of the pair (X_{L}, ∂X_{L}) maps ∂[F] = [∂F]. With respect to this framing, we can consider the Dehn filling of slope zero, resulting in the closed 3-manifold M_{L}. By definition, to obtain M_{L} attach a disc to each of the attaching curves, and then afterwards fill each of the resulting boundary spheres with a 3-ball.

Definition 5.1. The framing of the boundary tori of X_{L} constructed above is called the Seifert framing. The Seifert surgery on L is the 3-manifold M_{L} constructed above.

Remark 5.2. For links there is no reason for this framing to agree with the zero-framing of each individual component.

Collapsing the complement of a tubular neighbourhood of the Seifert surface F gives rise to map S^{3}∖νL → S^{1} = Bℤ, which extends to a map from the Seifert surgery ϕ: M_{L} → Bℤ. To see this in more detail, parametrise a regular neighbourhood of F as F × [−1, 1], with F as F × {0}. The intersection of this parametrised neighbourhood with each component of ∂F determines a parametrised subset S^{1} × [−1, 1] ⊂ S^{1} × S^{1} ⊆ ∂F. Extend this to a subset D^{2} × [−1, 1] ⊂ D^{2} × S^{1} for each of the Dehn filling solid tori D^{2} × S^{1} in M_{L}. Now define

ϕ: M_{L} → S^{1} = Bℤ
x ↦ \begin{cases}e^{πit} & x = (f, t) ∈ (F ∪ ⨆^{m} D^{2}) × [−1, 1] \\ −1 & otherwise. \end{cases}

The map ϕ classifies the image of the fundamental class of the capped-off Seifert surface in M_{L}, in the sense that [ϕ] maps to [F ∪ ⨆^{m} D^{2}] under [M_{L}, S^{1}] ⟶^{≅} H^{1}(M_{L}; ℤ) ⟶^{≅} H_{2}(M_{L}; ℤ). Recall that the homology class [F ∪ ⨆^{m} D^{2}] ∈ H_{2}(M_{L}; ℤ) only depends on the isotopy class of L and so also the homotopy class of ϕ does not depend on the Seifert surface F. The manifold M_{L} together with the map ϕ defines an element [(M_{L}, ϕ)] ∈ Ω_{3}(Bℤ), where Ω_{k}(X) denotes the bordism group of oriented, topological k-dimensional manifolds with a map to X. Recall that cobordism is a generalised homology theory fulfilling the suspension axiom, see e.g. [tD08, Chapter 21] and [May99, Section 14.4]. As a consequence, we obtain

Ω̃_{3}(Bℤ) = Ω̃_{3}(S^{1}) = Ω̃_{3}(ΣS^{0}) ≅ Ω̃_{2}(S^{0}) = Ω_{2}(pt) = 0.

Thus Ω_{3}(Bℤ) ≅ Ω_{3}(pt) = 0 [Roh53]. The group Ω_{3}(Bℤ) ≅ Ω_{3} ⊕ Ω_{2} = 0 ⊕ 0 = 0 is trivial, and we can make use of this fact to define a signature defect invariant, as follows. For any oriented 3-manifold M with a map ϕ: M → Bℤ, we will define an integer for each complex number z ∈ S^{1}. Since Ω_{3}(Bℤ) = 0, there exists a 4-manifold W with boundary M and a map Φ: W → Bℤ extending the map M → Bℤ on the boundary. Similarly to before, an element z ∈ S^{1} determines a representation

α: ℤ[π_{1}(W)] ⟶^{Φ} ℤ[ℤ] ⟶^{t↦z} ℂ.

Consider the twisted homology H_{i}(W; ℂ^{α}), and consider the intersection form λ_{α}(W) on the quotient H_{2}(W; ℂ^{α})/ im H_{2}(M; ℂ^{α}). Define the promised integer

σ(M, ϕ, z) := σ(λ_{α}(W)) − σ(W),

where σ(W) is the ordinary signature of the intersection form on W. The proof of the following proposition is known for the coefficient system ℚ(t), e.g. [Pow16]. For the convenience of the reader, we sketch the key steps for an adaptation to ℂ^{α}. Proposition 5.3.

(i) The intersection form λ_{α}(W) is nonsingular. (ii) The signature defect σ(M, ϕ, z) is independent of the choice of 4-

manifold W.

Proof. The long exact sequence of the pair (W, ∂W) = (W, M) gives rise to the following commutative diagram

\xymatrix{… \ar[r] & H_{2}(∂W; ℂ^{α}) \ar[r] & H_{2}(W; ℂ^{α}) \ar[r] \ar[ddr] & H_{2}(W, ∂W; ℂ^{α}) \ar[d]^{_{PD{−1 \atop W}}} \ar[r] & … \\  &  &  & H^{2}(W; ℂ^{α}) \ar[d]^{κ} \\  &  &  & (H_{2}(W; ℂ^{α}))^{∨} ,}

where for a ℂ-module P we denote its dual module by P^{∨} := Hom_{ℂ}(P, ℂ). Since Poincaré-Lefschetz duality PD_{W} and the Kronecker pairing κ are isomorphisms, we obtain an injective map H_{2}(W; ℂ^{α})/ im H_{2}(M; ℂ^{α}) → H_{2}(W; ℂ^{α})^{∨}. This map descends to

λ_{α}: H_{2}(W; ℂ^{α})/ im H_{2}(M; ℂ^{α}) → (H_{2}(W; ℂ^{α})/ im H_{2}(∂W; ℂ^{α}))^{∨} ,

so that the diagram below commutes:

\xymatrix{H_{2}(W; ℂ^{α})/ im H_{2}(M; ℂ^{α}) \ar[dr]_{λ_{α}} \ar@{^{(}->}[r] & H_{2}(W; ℂ^{α})^{∨} \\ & (H_{2}(W; ℂ^{α})/ im H_{2}(∂W; ℂ^{α}))^{∨} . \ar[u]} .

Consequently, the form λ_{α} is nondegenerate, and so it is nonsingular since it is a form over the field ℂ. We proceed with the second statement of the proposition, namely independence of σ(M, ϕ, z) on the choice of W. Suppose that we are given two 4-manifolds W^{+}, W^{−}, both with boundary ∂W^{±} = M, and a map Φ^{±}: W^{±} → Bℤ extending ϕ: M → Bℤ. Temporarily, define the signature defects arising from the two choices to be

σ(W^{±}, Φ^{±}, z) := σ(λ_{α}(W^{±})) − σ(W^{±}).

We will show that σ(W^{+}, Φ^{+}, z) = σ(W^{−}, Φ^{−}, z), and thus that σ(M, ϕ, z) is a well-defined integer, so our original notation was justified. Glue W^{+} and \overline{W^{−}} together along M, to obtain a closed manifold U, together with a map Φ: U → Bℤ. By Novikov additivity, we learn that

σ_{z}(U, Φ) := σ(λ_{α}(U)) − σ(U) = σ(W^{+}, Φ^{+}, z) − σ(W^{−}, Φ^{−}, z).

This defect σ_{z}(U, Φ) can be promoted to a bordism invariant σ_{z}: Ω_{4}(Bℤ) → ℤ, see e.g. [Pow16, Proof of Lemma 3.2] and replace ℚ(t) coefficients with ℂ^{α} coefficients.

Claim. The map σ_{z}: Ω_{4}(Bℤ) → ℤ is the zero map.

Let U be a closed 4-manifold together with a map Φ: U → S^{1}, representing an element of Ω_{4}(Bℤ). By the axioms of generalised homology theories, we have

Ω̃_{4}(S^{1}) = Ω̃_{4}(ΣS^{0}) ≅ Ω̃_{3}(S^{0}) = Ω_{3}(pt) = 0.

Thus an inclusion pt → S^{1} induced an isomorphism Ω_{4}(pt) ⟶^{≅} Ω_{4}(S^{1}). So (U, Φ) is bordant over S^{1} to a 4-manifold U′ with a null-homotopic map Φ′ to S^{1}. In this case the local coefficient system ℂ^{α} is just the trivial representation ℂ. Consequently, we have λ_{α}(U′) = λ(U′), so σ_{z}(U′, Φ′) = 0. By bordism invariance, σ_{z}(U, Φ) = 0, which completes the proof of the claim. Now the independence of σ(M, Φ, z) on the choice of W follows from

0 = σ_{z}(U, Φ) = σ(W^{+}, Φ^{+}, z) − σ(W^{−}, Φ^{−}, z).
□

Now that we have constructed an invariant, we need to relate it to the Levine-Tristram signatures. Recall that L is an oriented link, that M_{L} is the Seifert surgery, and that we constructed a canonical map ϕ: M_{L} → S^{1}, well-defined up to homotopy.

Lemma 5.4. Suppose that z ∈ S^{1}∖{1} and let ϕ: M_{L} → S^{1} be the map defined at the beginning of this section. Then we have

σ(M_{L}, ϕ, z) = σ_{L}(z).

Proof. Construct a 4-manifold with boundary M_{L} as follows. Let F be a connected Seifert surface for L. Push the Seifert surface into D^{4} and consider its complement V_{F} := D^{4}∖νF. Note that if we cap F off with m 2-discs, we obtain a closed surface. Let H be a 3-dimensional handlebody whose boundary is this surface. Note that ∂V_{F} = X_{L} ∪ F × S^{1}. Then define

W_{F} := V_{F} ∪_{F ×S^{1}} H × S^{1}.

Note that ∂W_{F} = M_{L}. By [Ko89, pp. 538-9] and [COT04, Lemma 5.4], we have that σ(W_{F}) = 0 and λ_{z}(W_{F}) = (1 − z)V + (1 − \overline{z})V^{T}. Therefore

σ(λ_{z}(W_{F})) − σ(W_{F}) = σ((1 − z)V + (1 − \overline{z})V^{T}) = σ(B(z)).
□

6. Concordance invariance of the signature

We start with a straightforward lemma, then we prove the final part of the main theorem. Recall that the complement X_{L} and the Seifert surgery M_{L} are both equipped with a homotopy class of a map to S^{1}, or equivalently with a cohomology class. For the link complement X_{L}, this class ξ_{L} ∈ H^{1}(X_{L}; ℤ) is characterised by the property that it sends each oriented meridian to 1.

Lemma 6.1. Let L and J be concordant links. Their Seifert surgeries M_{L} and M_{J} are homology bordant over S^{1}.

Proof. Denote the maps to S^{1} by ϕ_{L}: M_{L} → S^{1} and ϕ_{J}: M_{L} → S^{1}, and denote the corresponding cohomology classes by ξ_{L} ∈ H^{1}(M_{L}; ℤ) and ξ_{J} ∈ H^{1}(M_{J}; ℤ). Define X_{L} := S^{3}∖νL and X_{J} := S^{3}∖νJ. Let A ⊂ S^{3} × I be an embedding of a disjoint union of annuli giving a concordance between L and J. Fix a tubular neighbourhood νA = A × D^{2} of the annulus A with a trivialisation. Denote W_{A} := S^{3} × I∖νA, whose boundary consists of the union of X_{L}, X_{J}, and a piece identified with the total space of the unit sphere bundle A×S^{1} of νA. As usual, we refer to a representative {pt} × S^{1} for the S^{1} factor in A × S^{1} as a meridian of A. Note that the inclusions X_{L} ⊂ W_{A} and X_{J} ⊂ W_{A} map the meridians in the link complements to the meridians in W_{A}.

Claim. There exists a cohomology class ξ_{A} ∈ H^{1}(W_{A}; ℤ) mapping each meridian μ_{A} of A to 1.

This can be seen by the Mayer-Vietoris sequence

H^{1}(νA; ℤ) ⊕ H^{1}(W_{A}; ℤ) → H^{1}(∂νA; ℤ) → H^{2}(S^{3} × I; ℤ) = 0,

in which the map H^{1}(νA; ℤ) ≅ ℤ^{m} → H^{1}(∂νA; ℤ) ≅ (ℤ ⊕ ℤ)^{m} is given by 1 ↦ (1, 0) on each of the m summands. That is, the homology classes of the meridians of ∂νA ≅ A × S^{1} do not lie in the image of this surjective map, so they must lie in the image of H^{1}(W_{A}; ℤ). This completes the proof of the claim. It follows that ξ_{A} is pulled back to the unique classes ξ_{L} and ξ_{J} that map the meridians in the link complements to 1. Using the natural isomorphism between the functors [−, S^{1}] and H^{1}(−; ℤ), find a map ϕ_{W} : W_{A} → S^{1} that restricts to the prescribed map ϕ_{L} ⊔ ϕ_{J}: X_{L} ⊔ X_{J} → S^{1} on the boundary. Up to isotopy, there is a unique product structure on an annulus A = S^{1} × I. Having fixed such a structure, we consider the manifold

Y := W_{A} ∪_{A×S^{1}} ⨆^{m}(D^{2} × S^{1} × I).

The gluing is done in such a way as to restrict on ⨆^{m} S^{1} ×S^{1} ×{i}, for i = 0, 1, to the gluing of the Seifert surgery on X_{L} and X_{J}. By construction, this gives a bordism between M_{L} and M_{J}. Note that the map ϕ_{W} and the projection A × S^{1} → S^{1} glue together to give a map ϕ_{Y} : Y → S^{1}. Equipped with this map, (Y, ϕ_{Y}) is an S^{1}-bordism between (M_{L}, ϕ_{L}) and (M_{J}, ϕ_{J}).

Finally, we assert that Y is a homology bordism. To see this, first observe, as in the proof of Theorem 4.2, that W_{A} is a homology bordism from X_{L} to X_{J}. Flagrantly, A × S^{1} is a homology bordism from S^{1} × S^{1} to itself, and ⨆^{m}(D^{2} × S^{1} × I) is a homology bordism from ⨆^{m} D^{2} × S^{1} to itself. Gluing two homology bordisms together along a homology bordism, with the same maps on homology induced by the gluings for M_{L}, M_{J} and Y, it follows easily from the Mayer-Vietoris sequence and the five lemma that Y is a homology bordism. □

Theorem 6.2. Suppose that oriented m-component links L and J are concordant and that z ∈ S^{1}∖{1} is not a Knotennullstelle. Then σ_{L}(z) = σ_{J}(z).

Proof. As in the statement of the theorem, suppose that z ∈ S^{1}∖{1} is not a Knotennullstelle. Let W_{LJ} be a homology bordism between the Seifert surgeries M_{L} and M_{J}, whose existence is guaranteed by Lemma 6.1. Let W_{J} be a 4manifold that gives a nullbordism of M_{J} over Bℤ, and define W_{L} := W_{LJ} ∪_{M_{J}} W_{J}. The signature of the intersection form on H_{2}(W_{L}; ℂ^{a})/H_{2}(M_{L}; ℂ^{α}), together with the ordinary signature over ℤ, determines the signature σ_{L}(z) by Section 5. Similarly, the signature of the intersection form on the quotient H_{2}(W_{J}; ℂ^{α})/H_{2}(M_{J}; ℂ^{α}) and the ordinary signature of W_{J} determine the signature σ_{J}(z). By Lemma 3.1, we have homology isomorphisms

H_{2}(M_{L}; ℂ^{α}) ⟶^{≅} H_{2}(W_{LJ}; ℂ^{α}) and H_{2}(M_{J}; ℂ^{α}) ⟶^{≅} H_{2}(W_{LJ}; ℂ^{α}).

It follows that every class in H_{2}(W_{L}; ℂ^{α}) has a representative in W_{J}, that

H_{2}(W_{L}; ℂ^{a})/H_{2}(M_{L}; ℂ^{α}) ≅ H_{2}(W_{J}; ℂ^{a})/H_{2}(M_{J}; ℂ^{α}),

and that this isomorphism induces an isometry of the intersection forms. Thus the twisted signatures of both intersection forms are equal. We needed that z is not a Knotennullstelle in order to apply Lemma 3.1 in the preceding argument. The same argument over ℤ implies that the ordinary signatures also coincide, that is σ(W_{L}) = σ(W_{J}). Therefore σ(M_{L}, ϕ_{L}, z) = σ(M_{J}, ϕ_{J}, z), and so σ_{L}(z) = σ_{J}(z) by Lemma 5.4. Thus the Levine-Tristram signature at z is a concordance invariant, as desired. □

References

[Bai77] James L. Bailey. Alexander invariants of links. Thesis – The University of British Columbia. ProQuest LLC, Ann Arbor, MI, 1977.

[BZ03] Gerhard Burde and Heiner Zieschang. Knots, volume 5 of de Gruyter Studies in Mathematics. Walter de Gruyter & Co., Berlin, second edition, 2003.

[CF08] David Cimasoni and Vincent Florens. Generalized Seifert surfaces and signatures of colored links. Trans. Amer. Math. Soc., 360(3):1223– 1264 (electronic), 2008.

[CL04] Jae Choon Cha and Charles Livingston. Knot signature functions are independent. Proc. Amer. Math. Soc., 132(9):2809–2816 (electronic), 2004.

[COT03] Tim D. Cochran, Kent E. Orr, and Peter Teichner. Knot concordance, Whitney towers and L^{2}-signatures. Ann. of Math. (2), 157(2):433– 519, 2003.

[COT04] Tim D. Cochran, Kent E. Orr, and Peter Teichner. Structure in the classical knot concordance group. Comment. Math. Helv., 79(1):105– 123, 2004.

[FP12] Stefan Friedl and Mark Powell. An injectivity theorem for Casson-Gordon type representations relating to the concordance of knots and links. Bull. Korean Math. Soc., 49(2):395–409, 2012.

[FP14] Stefan Friedl and Mark Powell. Cobordisms to weakly splittable links. Proc. Amer. Math. Soc., 142:703–712, 2014.

[GL15] Patrick M. Gilmer and Charles Livingston. Signature jumps and Alexander polynomials for links. To appear in Proc. Amer. Math. Soc., 2015. Preprint available at arXiv:1508.04394.

[Gor78] Cameron Gordon. Some aspects of classical knot theory. In Knot theory (Proc. Sem., Plans-sur-Bex, 1977), pages 1–60. Springer, Berlin, 1978.

[Hil12] Jonathan Hillman. Algebraic invariants of links, volume 52 of Series on Knots and Everything. World Scientific Publishing Co. Pte. Ltd., Hackensack, NJ, second edition, 2012.

[Kau78] Louis H. Kauffman. Signature of branched fibrations. In Jean-Claude Hausmann, editor, Knot Theory, volume 685 of Lecture Notes in Mathematics, pages 203–217. Springer Berlin Heidelberg, 1978.

[Ko89] Ki Hyoung Ko. A Seifert-matrix interpretation of Cappell and Shaneson’s approach to link cobordisms. Math. Proc. Cambridge Philos. Soc., 106(3):531–545, 1989.

[Lev69] Jerome P. Levine. Invariants of knot cobordism. Invent. Math. 8 (1969), 98–110; addendum, ibid., 8:355, 1969.

[Lev07] Jerome P. Levine. Concordance of boundary links. J. Knot Theory Ramifications, 16(9):1111–1120, 2007.

[Lic97] W. B. Raymond Lickorish. An introduction to knot theory, volume 175 of Graduate Texts in Mathematics. Springer-Verlag, New York, 1997.

[May99] J. Peter May. A concise course in algebraic topology. Chicago Lectures in Mathematics. University of Chicago Press, Chicago, IL, 1999.

[Mur65] Kunio Murasugi. On a certain numerical invariant of link types. Trans. Amer. Math. Soc., 117:387–422, 1965.

[Pow16] Mark Powell. The four genus of a link, Levine-Tristram signatures and satellites. To appear in Journal of Knot Theory and its Ramifications, Special issue in honour of Tim Cochran, 2016. Preprint available at arXiv:1605.06833.

[Roh53] Vladimir A. Rohlin. Intrinsic homologies. Doklady Akad. Nauk SSSR (N.S.), 89:789–792, 1953.

[tD08] Tammo tom Dieck. Algebraic topology. EMS Textbooks in Mathematics. European Mathematical Society (EMS), Zürich, 2008.

[Tri69] Andrew G. Tristram. Some cobordism invariants for links. Proc. Cambridge Philos. Soc., 66:251–264, 1969.

Matthias Nagel Département de Mathématiques Université du Québec à Montréal Canada nagel@cirget.ca Mark Powell Département de Mathématiques Université du Québec à Montréal Canada mark@cirget.ca 44

Documenta Mathematica 22 (2017)